# LieRE: Lie Rotational Positional Encodings

### v4

@@ Abstract
Transformer architectures depend on explicit position encodings to capture token positional information. Rotary Position Encoding (RoPE) has emerged as a popular choice in language models due to its efficient encoding of relative position information through key-query rotations. However, RoPE faces significant limitations beyond language processing: it is constrained to one-dimensional sequence data and, even with learnable phases, offers limited representational capacity. We address these challenges with Lie Relative Encodings (LieRE), which generalizes RoPE to high-dimensional rotation matrices by leveraging their Lie group structure. Through extensive evaluation on three image datasets across 2D and 3D classification tasks, LieRE achieves 1.5% improvement over state-of-the-art baselines on 2D tasks and 1% on 3D tasks, while demonstrating superior generalization to higher resolutions. Our implementation is computationally efficient, with results reproducible on 4 A100 GPUs in 30 minutes on CIFAR100. Our code is available at https://github.com/StanfordMIMI/LieRE.
@@ 1 Introduction
Figure 1: Overview of how LieRE encodes spatial information, where $N$ denotes the number of input dimensions, i.e. $N=3$ for the 3D image. $A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x_{i}\;y_{i}\;z_{i}])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the $i$th patch in the flattened input, where $j$ corresponds to a different patch. $d$ is the head dimension.
@@
Figure 2: Comparisons of basis matrix $A$ of our work, Rope-Mixed (Heo et al., 2024), VisionLlama (Chu et al., 2024). First row: learnable parameter (yellow), not learnable parameter (blue) in basis matrix. Second row: rotation matrices $\exp(Ax)$ shared across the stem (gray) or different for each layer and head (colorful). MHA := Multi-Head-Attention, FF := Feed Forward, L[index] := Layer index, RxH := Rotation Matrix per Head, $\textrm{LieRE}_{\theta}$ := the number of LieRE parameters.
@@ 5 Experiments
In order to isolate the effect of changing the position encoding, we use a plain standard transformer backbone and training regime [^1] https://github.com/kentaroy47/vision-transformers-cifar10 with comparable baseline performance on CIFAR100 and ImageNet with (Lee et al., 2022). We implement relative position encoding types including LieRE, RoPE-Mixed, Visionllama, and absolute position encodings. We use the standard backbone sizes of ViT-Tiny, ViT-B and ViT-L (Dosovitskiy et al., 2020). All experiments use RandAugment (Cubuk et al., 2020). We avoid using pre-trained weights in order to maximize the comparability of results between methods. In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters and use the same default hyperparameters for all experiments (Appendix B.1). We evaluate two versions of LieRE, distinguished by the basis matrix tile sizes of 64 and 8, referred to as LieRE64 and LieRE8, respectively. Notably, a tile size of 2 corresponds to RoPE-Mixed.
5.1 2D Classification
For 2D data we evaluate performance on the CIFAR-100 and ImageNet-1k image classification task (Krizhevsky et al., 2009; Deng et al., 2009). We further analyse different backbone models sizes and data ablation effects. We examine accuracy both CIFAR-100 and ImageNet-1k and compare to absolute (Dosovitskiy et al., 2020) position encoding and recent related work, RoPE-Mixed (Heo et al., 2024), VisionLlama embeddings (Chu et al., 2024). For 2D image classification tasks, we demonstrate that the LieRE-based transformer achieves a relative performance improvement in accuracy of 10.0% over standard absolute position encoding (Touvron et al., 2022), 7.3% over RoPE adaptation in VisionLlama (Chu et al., 2024), and 2.2% over RoPE-Mixed (Heo et al., 2024) on CIFAR-100, with similar accuracy trends observed on ImageNet (Deng et al., 2009) (Table 1, Table 2).
We vary the capacity of the transformer backbone with corresponding to ViT-Tiny, ViT-B and ViT-L, as proposed in (Dosovitskiy et al., 2020). Additionally, we analyze the impact of incorporating LieRE on performance across different model sizes on the CIFAR-100 dataset, as shown in Table 6, Figure 8. Our results demonstrate that LieRE8 slightly outperforms alternatives across all evaluated model sizes.
@@
Our experiments reveal a pattern in how learnable relative position encodings perform under limited data conditions. Specifically, when we varied the available training data from 20% to 90% of the CIFAR-100 dataset, both LieRE and RoPE-Mixed demonstrated significantly better data efficiency compared to traditional transformer approaches for 2D image classification (figure 4).
5.2 Synthetic Spatial Classification
@@
Recent observations indicate that even advanced models like ChatGPT-4 and Claude Sonnet 3.5 struggle with basic spatial reasoning tasks. To investigate whether these limitations stem from positional encoding mechanisms, we designed a synthetic image classification task (Shah et al., 2024). The task presents a $108\times 108$ pixel image containing a $9\times 9$ grid (81 cells). Within this grid, we randomly place eight arrows and six letters (A, B, C, D, E, and Y), ensuring that one arrow is placed in the direction of the base of the ”Y”. The objective is to identify the direction of this specific arrow. To introduce visual complexity, we include four spurious letters and seven additional arrows as distractors. Figure 5 illustrates an example of this task setup.
@@
We train the models on 800,000 examples and observe that they generally converge after the first 400,000 examples. The only exception to this is absolute position encodings, where all variants were trained on 800,000 examples with no sign of increases in performance.
@@
Table 3 outlines the performance of different positional encoding methods on a synthetic task using a base model (85M parameters) equivalent to the DeiT architecture. We evaluate the models across three different input resolutions ($108\times 108$, $168\times 168$, and $276\times 276$ pixels), revealing differences in scalability and effectiveness. While absolute positional encoding shows degraded performance as resolution increases (from 45.1% to 40.1%) and RoPE-Mixed demonstrates strong but degrading performance at higher resolutions (from 100% to 88.6%), both LieRE variants maintain near-perfect accuracy across all resolution scales, with LieRE64 achieving 100% accuracy consistently, suggesting superior resolution invariance of our Lie group-based approach. Qualitative analysis shows that absolute and visionllama position encoding attend less clearly to the ”Y” token than RoPE-Mixed and LieRE. Please refer to the appendix for attention map examples, Figure 12 and Figure 12.
@@
In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-Mixed and absolute encoding similar to the previous section (Arnab et al., 2021; Heo et al., 2024). For the 3D experiments, we examine video classification performance in the UCF101 dataset (Soomro et al., 2012). Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from (Tong et al., 2022). The full set of hyperparameters may be found in appendix B.1. We observe a relative accuracy improvement of the LieRE-based transformer of up to 15.1% compared to absolute position embeddings and at least 1.5% compared to RoPE-inspired position encodings (table 1).
5.4 LieRE Capacity
LieRE adds a small amount of capacity to the model ($580k$ parameters 0.68% for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is. To answer this question, we vary the density of the learned basis and the learnability between heads and layers.
We vary capacity by enforcing a block-diagonal structure on the skew-symmetric basis. Varying the block size allows us to approach a dense matrix (LieRE64 for our models). Recall that using $2\times 2$ blocks recovers exact commutativity and is identical to RoPE-Mixed. As RoPE-Mixed is sensitive to the initialization of the weights, we initialize all generator weights mirroring the original presentation of RoPE-Mixed (Heo et al., 2024). We found that RoPE-Mixed is more sensitive to the initialization method than LieRE as seen in appendix table 7. In figure 6 we evaluate accuracy versus block dimension (Appendix Table 10).
@@ 5.4 LieRE Capacity
Table 4 presents the impact of sharing LieRE parameters across attention heads and layers on CIFAR-100 classification performance for RoPE-Mixed, LieRE64 and LieRE8 . We evaluate whether learning separate positional encodings for each attention head and layer provides performance benefits. We observe that learning across heads and layers yields superior performance followed by learning across layers.
5.5 Patch shuffling
Shuffling patches and frames allows us to see how much the model is able to use the positional information in its inputs. A model whose architecture does not allow/encourage the use of positional information will converge to a representation similar in spirit to a bag-of-words, where the relative locations of pixels/voxels do not matter. A greater dropoff in accuracy during shuffling is indicative that the model more heavily utilizes positional information.
We evaluate models using the decline in accuracy when evaluating on shuffled patches. We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE models are more capable at using positional information. The complete results are displayed for CIFAR-100 and table for UCF101 (table 5).
@@ 5.5 Patch shuffling
[…]
@@ 5.6 Multi-resolution Classification
In this section we compare the ability of methods to generalize to image resolutions not seen during training. We evaluate two training recipes inspired by (Heo et al., 2024). The first recipe matches the rest of the paper and consists of training the models on images of size $224\times 224$ for 200 epochs. The second adds an additional fine-tuning step at size $256\times 256$ for 30 epochs. The full details can be found in appendix B.1. We evaluate the accuracy on the ImageNet validation set with varying inference resolutions. Specifically, we scale the input images to resolutions of $196\times 196$, $256\times 256$, $320\times 320$, $384\times 384$, and $448\times 448$ pixels per dimension, and present the resulting accuracies in figure 7. For position assignment, we adopt a sequential approach where token positions are scaled proportionally to the image dimensions. For example, doubling the length of an image in each dimension doubles the range of positional indices. This method outperforms rescaling positions to a fixed range, as demonstrated by superior results for both RoPE-Mixed and LieRE across the evaluated training recipes.
@@ 6 Conclusion
In this paper, we introduced Lie group Relative position Encodings (LieRE), a position encoding that can effectively encode relative position information for attention mechanisms across modalities and input dimensionalities. Through experiments on 2D image classification (CIFAR-100, ImageNet-1k) and 3D video classification (UCF101), we demonstrated that LieRE achieves better performance compared to existing positional encoding methods. Beyond improving accuracy, LieRE also exhibits data and compute efficiency. On CIFAR-100, LieRE requires 3.5 times less training computed to match the performance of the baseline model with absolute position encodings. Furthermore, LieRE can outperform the baseline trained on the full dataset while using only 70% of the training data, highlighting its data efficiency. The key advantages of LieRE include its simplicity, flexibility, and ease of adaptation to new modalities. By requiring no changes to the tokenizer other than outputting positions and no other code changes, LieRE may provide a unified and efficient approach for transformers to process and learn from various data modalities within a single architecture.
@@ 7 Limitations
While LieRE shows promising results across multiple modalities and input dimensionalities, there are a few limitations worth noting. Our method is specifically designed to modify the inner product, making it compatible with most attention schemes, including original attention and linear attention. However, this specificity may limit its applicability to other architectures, like convolutional neural networks, that do not rely on inner product-based attention mechanisms. Future work could explore adaptations of our method to a wider range of architectures. Furthermore, the current formulation of our method is designed to encode vector positions in $\mathbb{R}^{d}$. While this is sufficient for many applications, it may not be directly applicable to tasks that require encoding poses in $SE(3)$ like robotics. Further research may be necessary to adapt our method to effectively handle such representational requirements.
Despite these limitations, we believe that our work provides much-needed insight into how to improve model performance and reduce training costs by encoding relative position information across various modalities and input dimensionalities.
@@ Acknowledgements
We would like to thank Maksim Maydanskiy for helping us understand Lie groups and Lie group representations. We would like to thank Aradhana Sinha for suggesting the shuffling experiment and providing early feedback on the paper. Sophie Ostmeier was supported by the German Research Foundation (DFG), Walter-Benjamin fellowship (ID: 517316550). This project was supported by Google Cloud and Azure credits.
@@ B.1 Experimental Hyperparameters
The CIFAR experiments where trained on 8xL4 GPUs with 24GB of VRAM each and all took under 30 minutes to complete. The basis capacity scaling experiment was conducted using RTX6000 GPUs. The ImageNet experiments were trained on 8xL40 GPUs and all took less than 2 days and 5 hours of runtime including time lost due to preemption and resource sharing. We use a cosine learning rate schedule with an initial learning rate of $41E-$ 4 and train for 200 epochs. We use an effective batch size of $512$. We use a patch size of $4\times 4$ on the original $32\times 32$ image for CIFAR-100 and a patch size of $16\times 16$ on the randomly cropped and resized $224\times 224$ image. All vision experiments used RandAugment (Cubuk et al., 2020). We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$. The hyperparameters were tuned with RoPE-Mixed and selected before conducting the LieRE trainers as to ensure a fair comparison.
@@
The 3D classification experiments were conducted on either $1008\times A$ 100 40GB GPUs or $1004\times A$ 100 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size. Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $2\times 16\times 16$ on the randomly cropped and resized $8\times 224\times 224$ video/image. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$.

### v5

@@ Abstract
Transformer architectures rely on position encodings to model the spatial structure of input data. Rotary Position Encoding (RoPE) is a widely used method in language models that encodes relative positions through fixed, block-diagonal, rotation matrices applied to key-query interactions. We hypothesize that this inductive bias limits their RoPE’s effectiveness for modalities with high dimensional structure. Lie Relative Encodings (LieRE) introduce a principled generalization of RoPE, aimed at increasing the representational capacity of positional encodings in transformers. Instead of fixed 2D rotations, LieRE learns dense skew-symmetric matrices (Lie algebra elements), which are then differentiable mapped to form high-dimensional rotation matrices (Lie group elements). This results in richer, learnable, and continuous, encodings of both relative and absolute positional information. We demonstrate the effectiveness of LieRE on 2D and 3D vision tasks, showing that it generalizes well to higher input resolutions while maintaining computational efficiency. The code and checkpoints are publicly available at https://github.com/StanfordMIMI/LieRE.
@@ 1 Introduction
Figure 1: Overview of how LieRE encodes spatial information, where NN denotes the number of input dimensions, i.e. $N=3$ for the 3D image. AA is a learnable skew symmetric matrix and $R_{i}=\exp(A[x_{i}\;y_{i}\;z_{i}])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the $i$th patch in the flattened input, where $j$ corresponds to a different patch. $d$ is the head dimension.
@@
Figure 2: Comparisons of basis matrix $A$ of our work, Rope-Mixed (Heo et al., 2024), VisionLlama (Chu et al., 2024). First row: learnable parameter (yellow), not learnable parameter (blue) in basis matrix. Second row: rotation matrices $\exp(Ax)$ shared across the stem (gray) or different for each layer and head (colorful). MHA := Multi-Head-Attention, $F$ := Feed Forward, L[index] := Layer index, RxH := Rotation Matrix per Head, $\textrm{LieRE}_{\theta}$ := the number of LieRE parameters.
@@ 5 Experiments
We evaluate LieRE across a range of tasks to assess its impact on transformer performance in 2D and 3D vision, spatial reasoning, and resolution generalization. Our goal is to isolate the contribution of positional encodings by using consistent architectures, training procedures, and hyperparameters across all methods. To this end, we utilize a standard recipe and build on top of the vision transformers repository (Yoshioka, 2024) and verify that our baselines perform similarly to prior work (Lee et al., 2022). We avoid using pre-trained weights in order to help reproducibility and comparability of the results between methods. We provide confidence interval estimates using bootstrap (B=1000, $\alpha=0.05$). We evaluate two versions of LieRE, distinguished by the basis matrix block-diagonal sizes of 64 and 8, referred to as LieRE64 and LieRE8, respectively. Notably, a tile size of 2 corresponds to RoPE-Mixed.
5.1 2D Image Classification
We begin with CIFAR-100 and ImageNet-1k benchmarks to evaluate LieRE in 2D vision tasks. All models use ViT-based architectures trained from scratch with standard data augmentations (RandAugment). We compare LieRE to absolute positional encodings, RoPE-Mixed (Heo et al., 2024), and VisionLLaMA (Chu et al., 2024). Table 1 shows that LieRE outperforms all baselines. On CIFAR-100, LieRE8 achieves a statistically significant relative improvement in top-1 accuracy relatively of 10.0% over absolute encodings, 7.3% over VisionLLaMA, and 2.2% over RoPE-Mixed. Similar trends hold on ImageNet (Table  2).
We also investigate performance across model sizes (ViT-Tiny, Base, Large). As shown in Table 6 and Figure 8, LieRE8 consistently outperforms other methods.
@@
To evaluate robustness in low-data regimes, we perform a data ablation study. Figure 4 shows that LieRE variants and RoPE-Mixed maintain significantly higher accuracy than baselines when training on only 20–90% of the CIFAR-100 dataset. This highlights LieRE’s data efficiency.
5.2 Synthetic Spatial Reasoning Task
@@
Recent observations indicate that even advanced models like ChatGPT-4 and Claude Sonnet 3.5 struggle with basic spatial reasoning tasks. To investigate whether these limitations stem from positional encoding mechanisms, we designed a synthetic image classification task (Shah et al., 2024). The task presents a $108\times 108$ pixel image containing a $9\times 9$ grid (81 cells). Within this grid, we randomly place eight arrows and six letters (A, B, C, D, E, and Y), ensuring that one arrow is placed in the direction of the base of the ”Y”. The objective is to identify the direction of this specific arrow. To introduce visual complexity, we include four spurious letters and seven additional arrows as distractors. Figure 5 illustrates an example of this task setup. We train the models on 800,000 examples and observe that they generally converge after the first 400,000 examples.
@@
Table 3 outlines the performance of different positional encoding methods on a synthetic task using a base model (85M parameters) equivalent to the DeiT architecture. We evaluate the models across three different input resolutions ($108\times 108$, $168\times 168$, and $276\times 276$ pixels), revealing differences in scalability and effectiveness. While absolute positional encoding shows degraded performance as resolution increases (from 45.1% to 40.1%) and RoPE-Mixed demonstrates strong but degrading performance at higher resolutions (from 100% to 88.6%), both LieRE variants maintain near-perfect accuracy across all resolution scales, with LieRE64 achieving 100% accuracy consistently. Qualitative analysis shows that absolute and visionllama position encoding attend less clearly to the ”Y” token than RoPE-Mixed and LieRE. Please refer to the appendix for attention map examples, Figure 12 and Figure 12.
@@
To assess LieRE’s performance on 3D data, we use the UCF101 video classification benchmark (Soomro et al., 2012). All models use a ViT-style backbone with 3D patch tokenization, trained from scratch with no hyperparameter tuning and the dataloader from (Tong et al., 2022). The full set of hyperparameters may be found in appendix B.1. We observe a relative accuracy improvement of the LieRE-based transformer of up to 15.1% compared to absolute position embeddings and at least 1.5% compared to RoPE-inspired position encodings (table 1).
5.4 LieRE Capacity and Parameter Efficiency
LieRE introduces minimal overhead–—only 580k additional parameters (0.68% for ViT-B)–—yet offers a flexible mechanism for increasing representational capacity. To explore the impact of this marginal capacity, we vary the density of the skew-symmetric basis and examine parameter sharing across heads and layers.
We control capacity via imposing a block-diagonal structure on the basis matrices. Smaller blocks (e.g., $2\times 2$) replicate RoPE-Mixed, while larger blocks increase expressivity, with LieRE64 using a fully dense basis. We adopt RoPE-Mixed’s initialization for fair comparison. We observe that RoPE-Mixed is more sensitive to initialization than LieRE (Appendix Table 7).
Figure 6 shows performance as a function of block size. Both in 2D (CIFAR-100) and 3D (UCF101), accuracy improves with block size, peaking around $8\times 8$—–suggesting this is a sweet spot between capacity and regularization.
@@ 5.4 LieRE Capacity and Parameter Efficiency
We also assess the effect of parameter sharing on CIFAR100 (Table 4). Learning LieRE parameters independently per head and per layer yields the best results. Sharing across layers or heads reduces accuracy, but still outperforms RoPE-Mixed.
5.5 Patch shuffling: Measuring Positional Dependency
Shuffling patches and frames allows us to see how much the model is able to use the positional information in its inputs. A model whose architecture does not allow/encourage the use of positional information will converge to a representation similar in spirit to a bag-of-words, where the relative locations of pixels/voxels do not matter. A greater drop-off in accuracy during shuffling is indicative that the model more heavily utilizes positional information.
We evaluate models using the decline in accuracy when evaluating on shuffled patches. We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE models rely more on positional information as expected. The complete results are displayed for CIFAR-100 and table for UCF101 (table 5).
@@ 5.5 Patch shuffling: Measuring Positional Dependency
$\uparrow Shuffling\downarrow$ $\uparrow$ $\uparrow Shuffling\downarrow$ $\uparrow$
@@ 5.6 Multi-resolution Classification
In this section we compare the ability of methods to generalize to image resolutions not seen during training. We evaluate two training recipes inspired by (Heo et al., 2024).
The first recipe matches the rest of the paper and consists of training the models on images of size $224\times 224$ for 200 epochs. The second adds an additional fine-tuning step at size $256\times 256$ for 30 epochs. The full details can be found in appendix B.1.
We evaluate the accuracy on the ImageNet validation set with varying inference resolutions. Specifically, we scale the input images to resolutions of $196\times 196$, $256\times 256$, $320\times 320$, $384\times 384$, and $448\times 448$ pixels per dimension, and present the resulting accuracies in figure 7.
For position assignment, we adopt a sequential approach where token positions are scaled proportionally to the image dimensions. For example, doubling the length of an image in each dimension doubles the range of positional indices. This method outperforms rescaling positions to a fixed range, as demonstrated by superior results for both RoPE-Mixed and LieRE across the evaluated training recipes.
@@ 6 Conclusion
We proposed Lie group Relative position Encodings (LieRE), a positional encoding method that modifies the attention method via dense, learned, high-dimensional rotation matrices. As compared to the more widely used block-2D rotation matrices typically used to encode positions, dense rotation matrices can encode both relative and absolute positional information and allow a greater portion of the model’s learnable capacity be allocated to spatial reasoning. Experiments on 2D image classification (CIFAR-100, ImageNet-1k) and 3D video classification (UCF101) show that LieRE consistently outperforms existing positional encoding methods. Our analysis indicates that LieRE-based ViTs effectively leverage spatial reasoning capabilities unavailable to transformers using only absolute positional encodings. In addition to accuracy gains, LieRE offers notable data and compute efficiency. Its simplicity, flexibility, and strong capacity to learn spatial structure make it broadly applicable. With no tokenizer modifications beyond position output and no additional architectural changes, LieRE provides simple approach for controlling the amount of positional information into transformers.
@@ 7 Limitations
While LieRE shows promising results for 2D and 3D inputs, several limitations are worth noting. For 1D input, LieRE reduces to RoPE with learnable phases (proof in appendix A). Our method is designed to modify the inner product, making it compatible with most attention mechanisms, including standard softmax attention and linear attention. However, this may limit its applicability to other architectures—–such as convolutional neural networks—–that do not rely on the attention mechanism. Future work could adapt the method to a broader range of architectures. The current formulation encodes vector positions in $\mathbb{R}^{d}$. While sufficient for many applications, it may not directly apply to tasks that require pose encoding in $SE(3)$ (e.g., robotics). Lastly, in its current implementation, LieRE relies on the accuracy and numerical stability of the matrix exponential in PyTorch. Future work may explore more efficient and robust implementations or approximations of this operation. Despite these limitations, we believe our approach provides valuable insight into improving model performance and reducing training costs by encoding relative positional information across various input dimensionalities.
@@ Acknowledgements
We would like to thank Maksim Maydanskiy for helping us understand Lie groups and Lie group representations. We would like to thank Aradhana Sinha for suggesting the shuffling experiment and providing early feedback on the paper. Furthermore, lucidrains suggested adding the proof of that in the 1D setting, LieRE has identical representational capacity to RoPE up to the learnable basis.Sophie Ostmeier was supported by the German Research Foundation (DFG), Walter-Benjamin fellowship (ID: 517316550). This project was supported by Google Cloud and Azure credits. This work was also supported in part by the Medical Imaging and Data Resource Center (MIDRC), which is funded by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) under contract 75N92020C00021 and through the Advanced Research Projects Agency for Health (ARPA-H).
@@ B.1 Experimental Hyperparameters
The CIFAR experiments where trained on 8xL4 GPUs with 24GB of VRAM each and all took under 30 minutes to complete. The basis capacity scaling experiment was conducted using RTX6000 GPUs. The ImageNet experiments were trained on 8xL40 GPUs and all took less than 2 days and 5 hours of runtime including time lost due to preemption and resource sharing. We use a cosine learning rate schedule with an initial learning rate of 1E41E-4 and train for 200 epochs. We use an effective batch size of $512$. We use a patch size of $4\times 4$ on the original $32\times 32$ image for CIFAR-100 and a patch size of $16\times 16$ on the randomly cropped and resized $224\times 224$ image. All vision experiments used RandAugment (Cubuk et al., 2020). We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$. The hyperparameters were tuned with RoPE-Mixed and selected before conducting the LieRE trainers as to ensure a fair comparison.
@@
The 3D classification experiments were conducted on either $8\times A100$ 40GB GPUs or $4\times A100$ 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size. Similar to 2D classification, we use an initial learning rate of 1E41E-4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $2\times 16\times 16$ on the randomly cropped and resized $8\times 224\times 224$ video/image. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$.